# Piotr Berman Meiram Murzabulatov

### v1

@@ 4.2 Analysis of the Construction
    where Section 4.2 was used to obtain the first inequality, and the last inequality holds for sufficiently small constant $c$.
    It remains to analyze $\Pr[E\mid\overline{G}],$ that is, the probability of $E$, given that the interesting window is bad. Consider a bad window of level $i\in\{\ell,\dots,h\}$. It has at most $q$ bridge squares that contain a queried pixel. Consider one of such bridge squares. Recall that this bridge square has $a_{i}/2-1$ bridges. Number these bridges using integers $11,2,\dots,a_{i}/2-$ 1. Let $x_{k}$ denote the number of pixels of Q on the bridge number $k\in[a_{i}/2-1]$ of the bridge square. Then the probability that this bridge square has a revealing set for the window is

### v2

@@ 4.2 Analysis of the Construction
where (4.2) was used to obtain the first inequality, and the last inequality holds for sufficiently small constant $c$.
It remains to analyze $\Pr[E\mid\overline{G}],$ that is, the probability of $E$, given that the interesting window is bad. Consider a bad window of level $i\in\{\ell,\dots,h\}$. It has at most $q$ bridge squares that contain a queried pixel. Consider one of such bridge squares. Recall that this bridge square has $a_{i}/2-1$ bridges. Number these bridges using integers $11,2,\dots,a_{i}/2-$ 1. Let $x_{k}$ denote the number of pixels of Q on the bridge number $k\in[a_{i}/2-1]$ of the bridge square. Then the probability over the choice of disconnecting pixels that this bridge square has a revealing set for the window is